In a survey, among 300 randomly selected single women, 69 of them said that they "definitely want to get married." In the same survey, among 400 randomly selected single men, 120 of them gave that same response. Assume that these samples are independent. Use a 0.01 signifficant level and the P-value method to test the claim that the proportion of single women who definitely want to get married is different from the one of single men who definitely want to get married. What is the P-value?

Select one:

a. 0.97%

b. 1.95%

c. 3.90%

d. None of the other answers is true.

e. 7.80%
Answer #1

solution - Here n,=300 x1=69, p = 69/300 = 0.23 ng = 40o, x2 = 120 82=120/400 = 0.3 Do nit2 = 69+120 = 0.27 nitre 300 tuoo He